OpenStudy (anonymous):

simplify the rational expression 4x^2+28x+40/x^2-2x-8

OpenStudy (ranga):

Factor 4 out of the numerator first. Then factor the remaining quadratic equations. Factor the quadratic expression in the denominator.

OpenStudy (anonymous):

i dont know how

OpenStudy (ranga):

I will do the numerator, you do the denominator: 4x^2+28x+40 = 4(x^2 + 7x + 10) = 4[ x^2 + 5x + 2x + 10 ] = 4[ x(x + 5) + 2(x + 5) ] = 4(x + 2)(x + 5) Denominator is: x^2-2x-8. You factor it.

OpenStudy (anonymous):

i dont know how please help

OpenStudy (ranga):

x^2-2x-8 = x^2 - 4x + 2x - 8 x(x - 4) + 2(x - 4) = (x + 2)(x - 4) Put the numerator and denominator together: 4(x + 2)(x + 5) / [ (x + 2)(x - 4) ] The (x + 2) cancels out. 4(x + 5) / (x - 4)

OpenStudy (anonymous):

THANK YOU SOOO MUCH i really appreciate it been dancing all day i cant even think ballet principal

OpenStudy (anonymous):

can you do another one ?

OpenStudy (ranga):

You are welcome. I can try. But you should learn these by practicing because quadratic equations will keep popping up.

OpenStudy (anonymous):

adding rational expressions 2/x+3/2x+1/4

OpenStudy (anonymous):

ugh i know i will

OpenStudy (anonymous):

Divide the rational expression 16x^3/x^2-25 divided by 10x/2x+10?

OpenStudy (anonymous):

thats it !

OpenStudy (ranga):

Does the second problem look exactly like this or does the x in the second term occur next to the 2 in the denominator?\[\Large \frac{ 2 }{ x } + \frac{ 3 }{ 2 }x + \frac{ 1 }{ 4 }\]

OpenStudy (anonymous):

pwease hurry YES IT DOES

OpenStudy (anonymous):

???

OpenStudy (ranga):

(3x^2 + 2x + 4) / 2x

OpenStudy (anonymous):

the x goes next to the 2 in 3 over 2

OpenStudy (anonymous):

?????

OpenStudy (ranga):

It is too late. for the second one. You are changing it at the last minute. Third one: 16x^2 / {5(x-5)}

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

can you show me work like how u got htat answer ?

OpenStudy (ranga):

2) (x + 20)/4x

OpenStudy (anonymous):

what is that for ? ^^

OpenStudy (ranga):

16x^3/x^2-25 divided by 10x/2x+10? 16x^3 / (x^2 - 25) / 10x/(2x+10) = 16x^3/[(x + 5)(x - 5)] * 2(x + 5) / 10x = 32x^2 / [10(x-5)] = 16x^2 / 5(x - 5)

OpenStudy (anonymous):

thank you !

OpenStudy (ranga):

You changed the problem and said the x in the second term of the second problem is in the bottom with 2. So the answer changed.

OpenStudy (anonymous):

one more ?

OpenStudy (anonymous):

ok so that was the answer

OpenStudy (ranga):

sorry. If you are doing an online test or something I shouldn't be giving you answers without you learning.